OpenStudy (justuu):

Medal Which pair of times has the longest elapsed time between the two times? 3:30 to 6:15 9:15 a.m. to 12:35 p.m. 6:45 to 10:00 11:10 p.m. to 2:15 a.m.

OpenStudy (unklerhaukus):

6:15 - 3:30 = 3:15 - 0:30 = 2:45

OpenStudy (unklerhaukus):

12:35 p.m. - 9:15 a.m. = 3:35 - 0:15 = 10:00 - 6:45 = 2:15a.m - 11:10 p.m. =

OpenStudy (justuu):

D? and thank you